An airplane takes 3 hours to travel 2160 miles with the wind. the return trip takes 4 hours against the wind. Find the speed of the airplane and speed of the wind.
I am unsure of how to approach this question and your assistance, or any advice would be much appreciated. Thank you.

Answer by richwmiller(17219)

rt=d
rate * time =distance
rw rate with the wind
rw*3=2160
solve for rw
rw=720
ra=rate against the wind
ra*4=2160
solve for ra
ra=540
rw-ra=wind/2
720-540=180/2=90
So the plane should fly at about 630 without wind.
one way the wind helped 630+90=720
the return trip the wind hindered 630-90=540
FYI Commercial jets actually fly in the range 510 to 565 mile
http://wiki.answers.com/Q/Speed_of_jet_planes_in_mph